OpenStudy (anonymous):

hey guys does anyone have any notes on differential calculus 1 or any useful tips or advice or does anyone have any challenging questions on this topic that i can practice? cause i have a big test on it next week and i want to do my best:) please u guise:)

OpenStudy (anonymous):

I have this printed on my wall last time http://www.eeweb.com/tools/math-sheets/pdfs/calculus-derivatives-limits.pdf , well i'm not sure if it will help you, but it helped me :)

OpenStudy (campbell_st):

I've always found these notes to be handy http://tutorial.math.lamar.edu/

OpenStudy (kamille):

are you guys doing this stuff in high school?

OpenStudy (anonymous):

@Data_LG2 and @campbell_st thanks so much guys really appreciate it! and yes we're doing this stuff in high school, but here in ireland we call high school secondary school:) just thought i'd let you know ^^,

OpenStudy (anonymous):

@Data_LG2 what does cot x mean ?? i have never come across cot before..

OpenStudy (anonymous):

it's basically 1/tan if i'm not mistaken

OpenStudy (anonymous):

huh? how does one get 'cot' what is cot made up of?

TheSmartOne (thesmartone):

cot is cotangent...

OpenStudy (anonymous):

i still don't know what cot is ...

OpenStudy (anonymous):

cot x = 1/tan x --> it's the inverse of tangent csc x= 1/sin x sec x = 1/cos x

OpenStudy (campbell_st):

from right angled trig you should know tan(x) = opp/adj. cot is another trig ratio and it is the reciprocal of tan. \[\cot(x) = \frac{1}{\tan(x)} = \frac{adj}{opp}\]